### v1

@@ Abstract
Perspective distortion (PD) leads to substantial alterations in the shape, size, orientation, angles, and spatial relationships of visual elements in images. Accurately determining camera intrinsic and extrinsic parameters is challenging, making it hard to synthesize perspective distortion effectively. The current distortion correction methods involve removing distortion and learning vision tasks, thus making it a multi-step process, often compromising performance. Recent work leverages the Möbius transform for mitigating perspective distortions (MPD) to synthesize perspective distortions without estimating camera parameters. An essential downside of using the Möbius transform is that it requires tuning multiple interdependent and interrelated parameters and involving complex arithmetic operations, leading to substantial computational complexity. To address these challenges, we propose Log Conformal Maps (LCM), a method leveraging the logarithmic function to approximate perspective distortions with fewer parameters and reduced computational complexity. We provide a theoretical foundation complemented with experiments to demonstrate that LCM with fewer parameters approximates the MPD. We show that LCM integrates well with supervised and self-supervised representation learning, outperform standard models, and matches the state-of-the-art performance in mitigating perspective distortion over multiple benchmarks, namely Imagenet-PD, Imagenet-E, and Imagenet-X. Further LCM demonstrate seamless integration with person re-identification and improved the performance. Source code shall be released to foster the research.
@@ 1 Introduction
Perspective distortion (PD) is a common issue in real-world imagery, complicating the development of computer vision applications. It arises from factors like camera position, depth, focal length, lens aberrations, and rotation, which affect the projection of 3D scenes onto 2D surfaces [20]. Accurately estimating these parameters for PD correction is challenging, posing a significant barrier to robust computer vision (CV) methods. Earlier studies focused on distortion correction [18], [27], [21], which makes vision task learning multi-steps. Recent work MPD [5] provides insights on mitigating perspective distortion by synthetically mimicking the PD in the learning process using nonlinear and conformal Möbius [1] transform.
@@
Previous research on image registration [22] demonstrates the potential of conformal log-polar mapping. Another study highlights the utility of complex logarithmic views [2] for exploring small details within extensive visual geometry contexts. Focusing on non-linear and conformal properties of log transform, we proposed Log Conformal Mapping (LCM) to mitigate perspective distortion and match state-of-the-art robustness performance. LCM method in Figure 1 demonstrates the synthesis of four perspective distorted views (left, right, top, bottom) mentioned in MPD [5]. The main contributions of this paper are as follows:
  1. 1.
    To the best of our knowledge, this work is the first to employ Log Conformal Transform for synthesizing perspective distortion, introducing the Log Conformal Maps (LCM) method.
  2. 2.
    We provide a theoretical foundation demonstrating that LCM, with fewer parameters, can approximate the Möbius Transforms-based MPD method to mitigate perspective distortion.
  3. 3.
    We validate our approach through extensive experiments on multiple perspective distortion-affected benchmarks, showing improvements in the robustness of both supervised and self-supervised models and showing effectiveness on real-world application, person re-identification.
@@ 2 Related Work
Earlier works have advanced computer vision tasks under perspective distortion. [24] introduced a method to correct perspective distortions in human mesh reconstruction from images with varying focal lengths. [16] proposed a framework for estimating camera parameters in the wild to improve 3D human pose and shape estimation. However, these methods rely on an inefficient two-step process and do not focus on robust representation learning for computer vision tasks.
@@ 4.1.1 Supervised learning
From Figure 2, we observe that a fully supervised LCM adapted ImageNet trained model shows higher accuracy than standard ImageNet trained ResNet50 [19] and retains similar performance with highly compute-intensive MPD [5]. We also compare the LCM with other popular data augmentation methods in Table 2 on ImageNet-PD, where LCM demonstrates a significant performance lead over other methods in supervised approach.
Figure 2: Comparison of LCM with MPD method [5] across probability values. Red line shows standard ResNet50. LCM outperform ResNet50 model trained on ImageNet and matches the performance with MPD. MPD results are reported from [5].
@@ 4.1.2 Self-supervised Representation learning
We investigate the robustness of LCM in the SimCLR self-supervised learning method, where pretraining is performed with LCM added as an augmentation with a probability of 0.8. Fine-tuning is then conducted following the protocol outlined in Section 4.1.1. Results for the self-supervised representation are provided in Table 3. As shown in Figure 3, LCM-trained ImageNet weights, when fully fine-tuned, outperform the standard ImageNet training and demonstrate competitive performance compared to the highly compute-intensive MPD method. Similar to the supervised approach, self-supervised trained LCM models also showcase significant performance improvement compared to other popular data augmentations (refer Table 2).
Table 3: LCM trained with SimCLR[4] self-supervised pretraining with probability $0.8$ and evaluated on ImageNet-PD subsets and original ImageNet. The probability $P$ of applying LCM in model fine-tuning stage. Results for probability $P$ denotes standard training and reported from MPD [5]. PD - Perspective Distorted.
@@
Figure 4: Comparison of linear evaluation performance on self-supervised trained models. (Left): Comparing original SimCLR [4], LCM+SimCLR and MPD+SimCLR[5] on original ImageNet and ImageNet-PD subsets. Performance on ImageNet-PD is average over all four subsets. (Right): Comparing original DINO [3], LCM+DINO and MPD+DINO [5] on original ImageNet and ImageNet-PD subsets. Performance on ImageNet-PD is average over all four subsets.
@@ 4.2 Qualitative Analysis
In this section, we present qualitative results in Figure 5 to assess the robustness of our proposed formulation. We have plotted the Grad-CAM for the LCM-trained SimCLR method and attention maps for the LCM-trained DINO method, comparing these with the Möbius and LCM methods. The Grad-CAM results for the LCM-trained SimCLR method show clearly defined attention regions, indicating effective feature localization. Similarly, the attention maps for the LCM-trained DINO method demonstrate precise focus areas. These visualizations highlight that LCM, with less computation and fewer parameters, produces qualitative outcomes comparable to the more compute-intensive MPD method [5].
@@ 4.3 ImageNet-E
ImageNet-Editing, commonly known as ImageNet-E [17], is a comprehensive dataset for benchmarking robustness against various object attributes, including perspective distortion. We extensively evaluated the LCM-trained models, both supervised and self-supervised, and compared their performance with other methods. The results are presented in Table 6 and Figure 6. In Table 6, we report both the drop in accuracy and the absolute accuracy for the fully supervised ResNet-50 network and the self-supervised SimCLR [4] method. The main observation is a significant boost in accuracy when LCM is incorporated into the training procedure, benefiting both background and object size changes in the dataset’s images. Figure 6 compares our proposed LCM-based technique with the existing Möbius-based MPD [5] method, demonstrating that the LCM-trained models outperform standard trained models overall the subsets and offers competitive performance compared with MPD [5] with fewer parameters and less compute-intensive, showing resilient for robustness.
Table 6: ImageNet-E: (a) Drop of Top1 accuracy (Lower is better) and (b) Absolute Accuracy (Higher is better) under background changes, size changes, random position (rp), random direction (rd) categories. Results for ResNet50 are reported from MPD [5].
(a) Drop of Accuracy
Original Background changes Size changes Position Direction
Models Inver $\lambda$ = -20 $\lambda$ = 20 $\lambda$ = 20-adv Random Full 0.10 0.08 0.05 rp rd
standard ResNet50 92.69 1.97 7.30 13.35 29.92 13.34 2.71 7.25 10.51 21.26 26.46 25.12
LCM (ResNet50) 92.60 1.72 6.70 11.55 29.95 13.64 3.20 6.08 8.7 18.54 23.58 22.92
LCM+SimCLR (ResNet50) 92.55 1.40 7.36 12.23 31.53 14.15 2.96 6.84 10.57 20.55 24.23 23.39
(b) Absolute Accuracy
Original Background changes Size changes Position Direction
Models Inver $\lambda$ = -20 $\lambda$ = 20 $\lambda$ = 20-adv Random Full 0.10 0.08 0.05 rp rd
standard ResNet50 92.69 90.72 85.39 79.34 62.77 79.35 89.98 85.44 82.18 71.43 66.23 67.57
LCM (ResNet50) 92.60 90.88 85.90 81.05 62.65 78.96 89.40 86.52 83.90 74.06 69.02 69.68
LCM+SimCLR (ResNet50) 92.55 91.15 85.19 80.32 61.02 78.40 89.59 85.71 81.98 72.00 68.32 69.16
@@ 4.4 ImageNet-X
ImageNet-X [14] contains human annotations identifying failure types in the widely-used ImageNet dataset. These annotations label distinguishing object factors such as pose, size, color, lighting, occlusions, and co-occurrences for each image in the validation set and a random subset of 12,000 training samples. We evaluated LCM-trained ImageNet models on ImageNet-X, and our findings are reported in Table 7 and Figure 7. From Table 7, it is evident that LCM-trained fully supervised ImageNet weights and self-supervised ImageNet weights offer better accuracies than the standard weights. Notably, LCM-trained weights demonstrate better or more competitive performance when compared to MPD-trained ImageNet weights. Figure 7 illustrates a reduction in the error ratio towards the optimal value of 1.0, decreasing from 1.55 to 1.43 in fully supervised experiments. A similar trend is observed for the self-supervised method, indicating increased robustness.
Table 7: ImageNet-X: error ratios comparisons for MPD and LCM. Error ration close to 1.0 is ideal robustness for each factor. LCM demonstrate consistent robustness compared to standard ResNet50 model and matches the performance with MPD in supervised and self-supervised approaches.
Factor Supervised Self-supervsied
Standard ResNet50 MPD LCM MPD+SimCLR LCM+SimCLR
pose 0.70 0.72 0.73 0.74 0.72
pattern 0.77 0.89 0.81 0.78 0.82
partial_view 1.13 0.89 0.86 0.86 0.87
background 1.14 1.08 1.10 1.09 1.11
brighter 1.17 1.09 1.10 1.09 1.09
color 1.19 1.21 1.17 1.18 1.16
larger 1.32 1.30 1.24 1.22 1.24
smaller 1.50 1.44 1.51 1.50 1.51
darker 1.72 1.52 1.50 1.51 1.52
shape 1.75 1.63 1.64 1.61 1.59
object_blocking 1.92 1.66 1.62 1.61 1.60
style 1.93 1.73 1.75 1.63 1.68
multiple_objects 1.97 1.76 1.72 1.72 1.69
subcategory 2.10 1.92 1.94 1.90 1.84
person_blocking 2.17 1.95 1.98 1.97 2.10
texture 2.39 2.20 2.24 2.18 2.20
Average 1.55 1.44 1.43 1.41 1.42
@@ 4.5 LCM Compute Efficiency
To demonstrate the computational efficiency of the proposed LCM method over the existing Möbius Transform-based MPD method [5], we performed a FLOP analysis, with the details presented in Table 8. Notably, LCM requires fewer floating point operations per second (FLOPs) than the MPD method. The reduced FLOPs are primarily due to LCM’s less complex arithmetic operations compared to Möbius transformations. Our benchmarking of the compute time for image transformation operations using LCM and MPD on a CPU-only setup further underscores the efficiency of the LCM method. The results show that MPD requires 0.24 seconds per image per transform, while LCM, with its streamlined approach, needs only 0.22 seconds per image per transform when using a single-core CPU.
@@
Table 9 shows the compute efficiency of LCM over MPD.
Table 9: Compute efficiency and power consumption comparison between MPD and LCM.
@@
Figure 8: LCM training time efficiency over probabilities range.
Figure 9: FLOPs comparison for LCM and MPD.
@@ 5 Person Re-identification
Following MPD [5], we integrated LCM with CLIP-ReIdent [10], a CLIP-based contrastive image-image pre-training for person re-identification on DeepSportRadar dataset [23] having video frames of varying poses and camera angles of players (Fig. 10).
Figure 10: LCM-Clip-ReIdent: LCM transforms the input gallery and query in the training of Clip-ReIdent for person re-identification. LCM applied with a probability of 0.1.
Table 10: LCM improves Clip-ReIdent model performance by enabling robust representation through controlled synthesis of PD.
@@
LCM-Clip-ReIdent models outperform original Clip-ReIdent and improve mean-average precision (mAP) across backbone ViT-L-14 and ResNet50x16 (Table 10) for both approaches, with and without re-ranking.
@@ 6 Conclusion and Future Work
In this paper, we introduce the Log Conformal Maps (LCM) based transform to synthesize perspective distortion efficiently, enhancing model robustness against PD. Our theoretical analysis, supported by detailed experimental results, demonstrates LCM’s capability to model perspective distortion effectively. While efficient, LCM-adapted models in supervised and self-supervised approaches outperform the standard models and match the state-of-the-art performance compared to recent methods. Perspective distortion is prevalent in multiple computer vision tasks and real-world applications such as crowd counting, object detection, and person re-identification. Therefore, potential future directions include applying LCM in various use cases and investigating its generalizability to downstream computer vision tasks. Furthermore, we conducted additional analysis and experiments to showcase the computational efficiency of the proposed technique over existing methods. We hope this work motivates further research into using logarithmic conformal mapping to address geometrical challenges in computer vision, advancing the concept and its applications.

### v2

@@ Abstract
Perspective distortion (PD) leads to substantial alterations in the shape, size, orientation, angles, and spatial relationships of visual elements in images. Accurately determining camera intrinsic and extrinsic parameters is challenging, making it hard to synthesize perspective distortion effectively. The current distortion correction methods involve removing distortion and learning vision tasks, thus making it a multi-step process, often compromising performance. Recent work leverages the Möbius transform for mitigating perspective distortions (MPD) to synthesize perspective distortions without estimating camera parameters. Möbius transform requires tuning multiple interdependent and interrelated parameters and involving complex arithmetic operations, leading to substantial computational complexity. To address these challenges, we propose Log Conformal Maps (LCM), a method leveraging the logarithmic function to approximate perspective distortions with fewer parameters and reduced computational complexity. We provide a detailed foundation complemented with experiments to demonstrate that LCM with fewer parameters approximates the MPD. We show that LCM integrates well with supervised and self-supervised representation learning, outperform standard models, and matches the state-of-the-art performance in mitigating perspective distortion over multiple benchmarks, namely Imagenet-PD, Imagenet-E, and Imagenet-X. Further LCM demonstrate seamless integration with person re-identification and improved the performance. Source code is made publicly available at https://github.com/meenakshi23/Log-Conformal-Maps.
@@ 1 Introduction
Perspective distortion (PD) is a common issue in real-world imagery, complicating the development of computer vision applications. It arises from factors like camera position, depth, focal length, lens aberrations, and rotation, which affect the projection of 3D scenes onto 2D surfaces [20]. Accurately estimating these parameters for PD correction is challenging, posing a significant barrier to robust computer vision (CV) methods.
@@
Earlier studies focused on distortion correction [18], [27], [21], which makes vision task learning multi-steps. Recent work MPD [5] provides insights on mitigating perspective distortion by synthetically mimicking the PD in the learning process using nonlinear and conformal Möbius [1] transform.
Previous research on image registration [22] demonstrates the potential of conformal log-polar mapping. Another study highlights the utility of complex logarithmic views [2] for exploring small details within extensive visual geometry contexts. Focusing on non-linear and conformal properties of log transform, we proposed Log Conformal Mapping (LCM) to mitigate perspective distortion and match state-of-the-art robustness performance. LCM method in Figure 1 demonstrates the synthesis of four perspective distorted views (left, right, top, bottom) mentioned in MPD [5]. The main contributions of this paper are:
1. To the best of our knowledge, this work is the first to employ Log Conformal Transform for synthesizing perspective distortion, introducing the Log Conformal Maps (LCM) method.
2. We present analytical and empirical evidence demonstrating that LCM, with fewer parameters, effectively mitigates perspective distortion, comparable to the Möbius Transforms-based MPD method.
2. We validate our approach through extensive experiments on multiple perspective distortion-affected benchmarks, showing improvements in the robustness of both supervised and self-supervised models and showing effectiveness on real-world application, person re-identification.
@@ 2 Related Work
Earlier works have advanced computer vision tasks under perspective distortion. [24] introduced a method to correct perspective distortions in human mesh reconstruction from images with varying focal lengths. [16] proposed a framework for estimating camera parameters in the wild to improve 3D human pose and shape estimation. However, these methods rely on an inefficient two-step process and do not focus on robust representations for computer vision tasks.
@@ 4.1.1 Supervised learning
From Figure 2, we observe that a fully supervised LCM adapted ImageNet trained model shows higher accuracy than standard ImageNet trained ResNet50 [19] and retains similar performance with highly compute-intensive MPD [5].
Figure 2: Comparison of LCM with MPD method [5] across probability values. Red line shows standard ResNet50. LCM outperform ResNet50 model trained on ImageNet and matches the performance with MPD. MPD results from [5].
We also compare the LCM with other popular data augmentation methods in Table 2 on ImageNet-PD, where LCM demonstrates a significant performance lead over other methods in supervised approach.
@@ 4.1.2 Self-supervised Representation learning
We investigate the robustness of LCM in the SimCLR self-supervised learning method, where pretraining is performed with LCM added as an augmentation with a probability of 0.8. Fine-tuning is then conducted following the protocol outlined in Section 4.1.1. Results for the self-supervised representation are provided in Table 3. As shown in Figure 3, LCM-trained ImageNet weights, when fully fine-tuned, outperform the standard ImageNet training and demonstrate competitive performance compared to the highly compute-intensive MPD method. Similar to the supervised approach, self-supervised trained LCM models also showcase significant performance improvement compared to other popular data augmentations (Table 2).
Table 3: LCM trained with SimCLR[4] self-supervised pretraining with probability $0.8$ and evaluated on ImageNet-PD subsets and original ImageNet. The probability $P$ of applying LCM in model fine-tuning stage. Results for probability $P$ denotes standard training from MPD [5]. PD - Perspective Distorted.
@@
Figure 4: Comparison of linear evaluation performance on self-supervised trained models. (Left): original SimCLR [4], LCM+SimCLR and MPD+SimCLR[5] on original ImageNet and ImageNet-PD subsets. Performance on ImageNet-PD is average over all four subsets. (Right): Comparing original DINO [3], LCM+DINO and MPD+DINO [5] on original ImageNet and ImageNet-PD subsets. Performance on ImageNet-PD is average over all subsets.
@@ 4.2 Qualitative Analysis
In this section, we present qualitative results in Figure 5 to assess the robustness of our proposed formulation. We have plotted the Grad-CAM for the LCM-trained SimCLR method and attention maps for the LCM-trained DINO method, comparing these with the Möbius and LCM methods. The Grad-CAM results for the LCM-trained SimCLR method show clearly defined attention regions, indicating effective feature localization. Similarly, the attention maps for the LCM-trained DINO method demonstrate precise focus areas. These visualizations highlight that LCM, with fewer parameters, produces comparative qualitative outcomes to MPD  [5].
@@ 4.3 ImageNet-E
ImageNet-Editing, commonly known as ImageNet-E [17], is a comprehensive dataset for benchmarking robustness against various object attributes, including perspective distortion. We extensively evaluated the LCM-trained models, both supervised and self-supervised, and compared their performance with other methods. The comparison are presented in Figure 6 and detailed results are in Table 9 in suppl. material. The main observation is a significant boost in accuracy when LCM is incorporated into the training procedure, benefiting both background and object size changes in the dataset’s images. Figure 6 compares our proposed LCM-based technique with the existing Möbius-based MPD [5] method, demonstrating that the LCM-trained models outperform standard trained models overall the subsets and offers competitive performance compared with MPD [5] with fewer parameters and less compute-intensive, showing resilient for robustness.
@@ 4.4 ImageNet-X
ImageNet-X [14] contains human annotations identifying failure types in the widely-used ImageNet dataset. These annotations label distinguishing object factors such as pose, size, color, lighting, occlusions, and co-occurrences for each image in the validation set and a random subset of 12,000 training samples. We evaluated LCM-trained ImageNet models on ImageNet-X, and our findings are reported in Figure 7 and detailed results are in Table 10 in suppl. material. Figure 7 illustrates a reduction in the error ratio towards the optimal value of 1.0, decreasing from 1.55 to 1.43 in fully supervised experiments. A similar trend is observed for the self-supervised method, indicating increased robustness.
@@ 4.5 LCM Compute Efficiency
To demonstrate the computational efficiency of the proposed LCM method over the existing Möbius Transform-based MPD method [5], we performed a FLOP analysis, with the details presented in Table 6. Notably, LCM requires fewer floating point operations per second (FLOPs) than the MPD method. The reduced FLOPs are primarily due to LCM’s less complex arithmetic operations compared to Möbius transformations. Our benchmarking of the compute time for image transformation operations using LCM and MPD on a CPU-only setup further underscores the efficiency of the LCM method. The results show that MPD requires 0.24 seconds per image per transform, while LCM, with its streamlined approach, needs only 0.22 seconds per image per transform when using a single-core CPU. Table 7 shows the compute efficiency of LCM over MPD.
@@
Table 7: Compute efficiency and power consumption between MPD and LCM.
@@
Figure 8: Comparison of LCM and MPD for (a) training time efficiency and (b) FLOPs across probability ranges.
@@ 5 Person Re-identification
Following MPD [5], we integrated LCM with CLIP-ReIdent [10], a CLIP-based contrastive image-image pre-training for person re-identification on DeepSportRadar dataset [23] having video frames of varying poses and camera angles of players (Fig. 9). LCM-Clip-ReIdent models outperform original Clip-ReIdent and improve mean-average precision (mAP) across backbone ViT-L-14 and ResNet50x16 (Table 8) for both approaches, with and without re-ranking.
Figure 9: LCM-Clip-ReIdent: LCM transforms the input gallery and query during training of Clip-ReIdent for person re-id. LCM probability is set to 0.1.
Table 8: LCM improves Clip-ReIdent model.
@@ 6 Conclusion and Future Work
In this paper, we present the Log Conformal Maps (LCM) transform to efficiently handle perspective distortion, enhancing model robustness. Our analysis, backed by experiments, shows LCM’s effectiveness in modeling perspective distortion. LCM-adapted models in both supervised and self-supervised settings surpass standard models and achieve state-of-the-art performance. Given its relevance to tasks like crowd counting and object detection, future directions include exploring LCM’s application to diverse computer vision tasks and its computational advantages. We aim to inspire further research into using logarithmic conformal mapping for addressing geometric challenges in computer vision.
Acknowledgment: The authors thank Sumit Rakesh, Luleå University of Technology, for his support with the Lotty Bruzelius cluster. We also thank the National Supercomputer Centre at Linköping University for the Berzelius supercomputing, supported by the Knut and Alice Wallenberg Foundation.